OpenStudy (anonymous):

solve by substitution x/2-y/3=5/6 x/5-y/4=29/10

sam (.sam.):

(x)/(2)-(y)/(3)=(5)/(6) -(y)/(3)=-(x)/(2)+(5)/(6) -y=(-3x+5)/(2) y=(-3x+5)/(2)*-1 y=(3x-5)/(2) ---------------------------------------- (x)/(5)-((3x-5)/(2))/(4)=(29)/(10) (x)/(5)-(3x-5)/(8)=(29)/(10) (x)/(5)*(8)/(8)-(3x-5)/(8)*(5)/(5)=(29)/(10) (x(8))/(40)-((3x-5)(5))/(40)=(29)/(10) (x(8)-((3x-5)(5)))/(40)=(29)/(10) (-7x+25)/(40)=(29)/(10) -7x+25=116 -7x=91 x=-13 -------------------------------------- y=-22 =========================== (-13,-22)

OpenStudy (anonymous):

wow thanks sososososososo mucho